For any linear equation why is it true that for the x intercept, the y coordinate is zero and for the y intercept the coordinate is zero?
Found 2 solutions by Alan3354, josgarithmetic:
Answer by Alan3354(69443)

For any linear equation why is it true that for the x intercept, the y coordinate is zero and for the y intercept the coordinate is zero?
-----------
The y of any point is the distance from the x-axis. At the x-int, the distance to the x-axis = 0 --> (x,0) --> y = 0.
---
And vice versa.

Answer by josgarithmetic(39838)

An axis intercept is the coordinate on the axis which intersects with or shares a point on the line. This then is the description about which you asked.

Ax+By=C standard form for a line equation.

Let x=0. What is y?
A*0+By=C
0+By=C
y=C/B
The y-intercept is (0, C/B)

Let y=0: What is x?
Ax+B*0=C
Ax=C
x=C/A
The x-intercept is (C/A, 0)